# Agentic reinforcement learning empowers next-generation chemical language models for molecular design and synthesis

### v1

@@ Abstract
Language models are revolutionizing the biochemistry domain, assisting scientists in drug design and chemical synthesis with high efficiency. Yet current approaches struggle between small language models prone to hallucination and limited knowledge retention, and large cloud-based language models plagued by privacy risks and high inference costs. To bridge this gap, we introduce ChemCRAFT, a novel framework leveraging agentic reinforcement learning to decouple chemical reasoning from knowledge storage. Instead of forcing the model to memorize vast chemical data, our approach empowers the language model to interact with a sandbox for precise information retrieval. This externalization of knowledge allows a locally deployable small model to achieve superior performance with minimal inference costs. To enable small language models for agent-calling ability, we build an agentic trajectory construction pipeline and a comprehensive chemical-agent sandbox. Based on sandbox interactions, we constructed ChemToolDataset, the first large-scale chemical tool trajectory dataset. Simultaneously, we propose SMILES-GRPO to build a dense chemical reward function, promoting the model’s ability to call chemical agents. Evaluations across diverse aspects of drug design show that ChemCRAFT outperforms current cloud-based LLMs in molecular structure analysis, molecular optimization, and synthesis pathway prediction, demonstrating that scientific reasoning is not solely an emergent ability of model scale, but a learnable policy of tool orchestration. This work establishes a cost-effective and privacy-preserving paradigm for AI-aided chemistry, opening new avenues for accelerating molecular discovery with locally deployable agents.
@@ Introduction
Chemical Language Models (CLMs) have emerged as transformative tools in accelerating drug discovery and materials science [1, 2, 3], demonstrating remarkable potential in tasks ranging from molecular property prediction [4, 5, 6] and de novo design [7, 8, 9] to retrosynthesis planning [10, 11, 12] and reaction condition recommendation [13, 14, 15]. Currently, the development of CLMs primarily follows two distinct paradigms. The first approach involves supervised fine-tuning (SFT), or the continued pre-training of large language models on domain-specific corpora [16, 17, 18, 19, 20]. While these models achieve competitive performance on in-distribution metrics [17, 21, 22], they fundamentally suffer from rigid optimization objectives. By forcing the model to minimize loss on "Task-Query-Direct Answer" triplets, current SFT paradigms compel the LLM to learn a shallow mapping from chemical inputs directly to numerical values or labels [23, 24, 25]. Critically, this approach bypasses the "expert-like" reasoning process—such as structural analysis, intermediate hypothesis generation, and logical verification—that is essential for scientific discovery [26, 27, 28]. Consequently, this leads to a degradation of the model’s intrinsic general capabilities, often referred to as catastrophic forgetting [29, 30, 31], and limits its adaptability across disparate chemical tasks [32, 33, 34].
@@
To rigorously validate the application potential of our framework across diverse chemical domains, we employ ChemCoTBench, a comprehensive evaluation suite designed to assess the multi-step reasoning and tool-use capabilities of chemical agents. Covering a spectrum of nine critical tasks—ranging from fundamental molecular structure understanding and editing to molecular property optimization and reaction-related tasks—this benchmark provides a holistic view of an agent’s proficiency. Extensive evaluations reveal that ChemCRAFT achieves a new state-of-the-art among open-source models. Remarkably, despite its compact parameter size, it demonstrates problem-solving capabilities that not only significantly surpass the baselines of similar scales but also rival, and in specific reasoning-intensive tasks, exceed those of leading commercial LLM APIs.
@@ 1.2 Chemical Proficiency Evaluation
To evaluate chemical intuition at the systemic level, we assessed models on three pillars of reaction prediction: (1) Forward Prediction, encompassing both Major Products and, crucially, By-Products (vital for assessing impurity profiles and purification risks); (2) Retrosynthesis (w/o reaction type), testing the ability to deconstruct targets into available precursors; and (3) Condition Recommendation, requiring precise suggestions for reagents, solvents, and catalysts. We benchmarked against domain-specialized models (e.g., Chemformer [64], GraphRetro [65], RetroPrime [66]) and general reasoning LLMs (e.g., GPT-o3, DeepSeek-R1).
@@
The model’s superior grasp of reactivity is further evidenced in the case studies (Figure 2e). We selected challenging modern reactions involving non-polar bond activation (C-C, C-H, C-F), transformations that are underrepresented in standard patent data (USPTO) and typically confound expert models. In the Ruthenium-catalyzed C-C coupling case (top), Gemini-2.5-Pro misses the competitive advantage of C-O activation over remote C-H activation. In the MnO2 oxidation case (middle), it underestimates the oxidation depth, predicting an amide rather than the correct carbonyl product. In the Cobalt-catalyzed C-F activation (bottom), the baseline fails to recognize that the specific additives are designed to suppress the benzyne pathway, leading to a regio-isomer error. ChemCRAFT correctly predicts these outcomes not by mere intuition, but by evidence-based reasoning: its sandbox retrieves analogous reaction templates from the library, allowing the agent to ground its mechanism in empirical precedent rather than hallucinating plausible but incorrect pathways.
@@ 2 Methods
We present a unified framework designed to empower small language models [68, 69]with autonomous reasoning and operational capabilities in the chemical domain. As illustrated in Fig.1, our methodology is orchestrated across three interconnected pillars. First, to overcome the limitations of static knowledge storage, we construct a Chemical Agent Sandbox, an interactive environment that externalizes domain expertise through computational, deep learning, and retrieval-based tools. Second, to bridge the gap between abstract chemical knowledge and actionable problem-solving, we introduce a pipeline for constructing High-Quality Reasoning Trajectories, incorporating tool-integrated narratives and reflective refinement. Finally, we implement a Two-Stage Training Paradigm, which progressively transitions the model from supervised behavioral cloning to robust reinforcement learning, optimizing for scientific validity and rigorous chemical constraints via the SMILES-GRPO mechanism.

### v2

@@ Abstract
Language models are revolutionizing the biochemistry domain, assisting scientists in drug design and chemical synthesis with high efficiency. Yet current approaches struggle between small language models prone to hallucination and limited knowledge retention, and large cloud-based language models plagued by privacy risks and high inference costs. To bridge this gap, we introduce ChemCRAFT, a novel framework leveraging agentic reinforcement learning to decouple chemical reasoning from knowledge storage. Instead of forcing the model to memorize vast chemical data, our approach empowers the language model to interact with a sandbox for precise information retrieval. This externalization of knowledge allows a locally deployable small model to achieve superior performance with minimal inference costs. To enable small language models for agent-calling ability, we build an agentic trajectory construction pipeline and a comprehensive chemical-agent sandbox. Based on sandbox interactions, we constructed ChemToolDataset, the first large-scale chemical tool trajectory dataset. Simultaneously, we propose SMILES-GRPO to build a dense chemical reward function, promoting the model’s ability to call chemical agents. Evaluations across diverse aspects of drug design show that ChemCRAFT outperforms current cloud-based LLMs in molecular structure analysis, molecular optimization, and synthesis pathway prediction, demonstrating that scientific reasoning is not solely an emergent ability of model scale, but a learnable policy of tool orchestration. This work establishes a cost-effective and privacy-preserving paradigm for AI-aided chemistry, opening new avenues for accelerating molecular discovery with locally deployable agents. Code available at https://github.com/HowardLi1984/ChemCraft.
@@ Introduction
Chemical Language Models (CLMs) have emerged as transformative tools in accelerating drug discovery and materials science [1, 2, 3, 4], demonstrating remarkable potential in tasks ranging from molecular property prediction [5, 6, 7] and de novo design [8, 9, 10, 11] to retrosynthesis planning [12, 13, 14] and reaction condition recommendation [15, 16, 17]. Currently, the development of CLMs primarily follows two distinct paradigms. The first approach involves supervised fine-tuning (SFT), or the continued pre-training of large language models on domain-specific corpora [18, 19, 20, 21, 22]. While these models achieve competitive performance on in-distribution metrics [19, 23, 24], they fundamentally suffer from rigid optimization objectives. By forcing the model to minimize loss on "Task-Query-Direct Answer" triplets, current SFT paradigms compel the LLM to learn a shallow mapping from chemical inputs directly to numerical values or labels [25, 26, 27]. Critically, this approach bypasses the "expert-like" reasoning process—such as structural analysis, intermediate hypothesis generation, and logical verification—that is essential for scientific discovery [28, 29, 30]. Consequently, this leads to a degradation of the model’s intrinsic general capabilities, often referred to as catastrophic forgetting [31, 32, 33], and limits its adaptability across disparate chemical tasks [34, 35, 36].
@@
To rigorously validate the application potential of our framework across diverse chemical domains, we employ ChemCoTBench [63], a comprehensive evaluation suite designed to assess the multi-step reasoning and tool-use capabilities of chemical agents. Covering a spectrum of nine critical tasks—ranging from fundamental molecular structure understanding and editing to molecular property optimization and reaction-related tasks—this benchmark provides a holistic view of an agent’s proficiency. Extensive evaluations reveal that ChemCRAFT achieves a new state-of-the-art among open-source models. Remarkably, despite its compact parameter size, it demonstrates problem-solving capabilities that not only significantly surpass the baselines of similar scales but also rival, and in specific reasoning-intensive tasks, exceed those of leading commercial LLM APIs.
@@ 1.2 Chemical Proficiency Evaluation
To evaluate chemical intuition at the systemic level, we assessed models on three pillars of reaction prediction: (1) Forward Prediction, encompassing both Major Products and, crucially, By-Products (vital for assessing impurity profiles and purification risks); (2) Retrosynthesis (w/o reaction type), testing the ability to deconstruct targets into available precursors; and (3) Condition Recommendation, requiring precise suggestions for reagents, solvents, and catalysts. We benchmarked against domain-specialized models (e.g., Chemformer [67], GraphRetro [68], RetroPrime [69]) and general reasoning LLMs (e.g., GPT-o3, DeepSeek-R1 [64]).
@@
The model’s superior grasp of reactivity is further evidenced in the case studies (Figure 2e). We selected challenging modern reactions involving non-polar bond activation (C-C, C-H, C-F), transformations that are underrepresented in standard patent data (USPTO) and typically confound expert models. In the Ruthenium-catalyzed C-C coupling case [70] (top), Gemini-2.5-Pro misses the competitive advantage of C-O activation over remote C-H activation. In the MnO2 oxidation case [71] (middle), it underestimates the oxidation depth, predicting an amide rather than the correct carbonyl product. In the Cobalt-catalyzed C-F activation [72] (bottom), the baseline fails to recognize that the specific additives are designed to suppress the benzyne pathway, leading to a regio-isomer error. ChemCRAFT correctly predicts these outcomes not by mere intuition, but by evidence-based reasoning: its sandbox retrieves analogous reaction templates from the library, allowing the agent to ground its mechanism in empirical precedent rather than hallucinating plausible but incorrect pathways.
@@ 2 Methods
We present a unified framework designed to empower small language models [74, 75, 76]with autonomous reasoning and operational capabilities in the chemical domain. As illustrated in Fig.1, our methodology is orchestrated across three interconnected pillars. First, to overcome the limitations of static knowledge storage, we construct a Chemical Agent Sandbox, an interactive environment that externalizes domain expertise through computational [77], deep learning, and retrieval-based [78, 79] tools. Second, to bridge the gap between abstract chemical knowledge and actionable problem-solving, we introduce a pipeline for constructing High-Quality Reasoning Trajectories, incorporating tool-integrated narratives and reflective refinement. Finally, we implement a Two-Stage Training Paradigm, which progressively transitions the model from supervised behavioral cloning to robust reinforcement learning, optimizing for scientific validity and rigorous chemical constraints via the SMILES-GRPO mechanism.